Use the following information to answer questions 16-18

Sales revenue

$778,000

Operating expenses excluding depreciation

Depreciation expense

Unrealized loss on FV-NI investments

Loss on sale of equipment

Income before income tax

Income tax expense

Net Income

$499,000

66,000

4,000

14,000

  

583,000

195,000

58,000

$137,000

There were no purchases or sales of trading (FV-NI) investments during 2014. Ellis’s statement of financial position included the following comparative data at December 31:

2014

             2013

Accounts receivable

$35,000

$54,000

FV-NI investments

22,000

26,000

Accounts payable

44,000

31,000

Income tax payable

6,000

8,500

1. How much cash was received from customers?

  1. $778,000
  2. $797,000
  3. $759,000
  4. $743,000

2. How much cash was paid for income taxes?

  1. $58,000
  2. $55,500
  3. $60,500
  4. $52,000

3. Assuming all accounts payable relate to operating expenses, how much cash was paid for operating expenses?

  1. $420,000
  2. $552,000
  3. $512,000
  4. $486,000

How to get these MCQS
Answer #1

Requirement 1

Cash received from customers = Sales Revenue + Accounts Receivable 2013 - Accounts Receivable 2014

                                             = 778000 + 54000 - 35000

                                             = 797000

Requirement 2

Cash paid for Income Taxes = Income Tax expense + Income Tax Payable 2013 - Income Tax Payable 2014

                                         = 58000 + 8500 - 6000

                                         = 60500

Requirement 3

Operating expenses paid = Operating expenses excluding depreciation + Accounts payable 2013 - Accounts payable 2014

                                     = 499000 + 31000 - 44000

                                     = 486000